# Many-body Fu-Kane-Mele index

### v1

@@ Abstract
We define a ${\mathbb{Z}}_{2}$-valued index for stably short-range entangled states of two-dimensional fermionic lattice systems with charge conservation and time reversal symmetry. The index takes its non-trivial value precisely if the ‘fluxon’, the state obtained by inserting a $\pi$-flux through the system, transforms under time reversal as part of a Kramers pair. This index extends the Fu-Kane-Mele index of free fermions to interacting systems.
@@ 1 Introduction
The invariant that characterises the quantum spin Hall effect is the ${\mathbb{Z}}_{2}$-valued Fu-Kane-Mele index [32, 31, 25, 26]. It was argued in [48] [47] that this invariant has a physical interpretation in terms of time reversal transformation properties of defects bound to ±$\pm\pi$ magnetic fluxes. If the invariant is non-trivial then these defects form a Kramers pair for the time reversal symmetry. We review the argument below. While these early works placed some importance on the role of spin, they all recognized that the invariant is in fact only relying on fermionic time reversal symmetry: no component of spin must be conserved for the invariant to be meaningful.
@@ 1.1 Idea behind the construction
The systems we consider are interacting generalisations of gapped free fermion Hamiltonians on the two-dimensional square lattice with a time reversal symmetry ${\mathscr{T}}$ that squares to $-\operatorname{\mathbbm{1}}$ (Altland-Zirnbauer class AII [1]). Such Hamiltonians come in two topological phases distinguished from each other by the ${\mathbb{Z}}_{2}$-valued Fu-Kane-Mele index. A nice characterisation of the index is obtained by piercing magnetic flux through the system [17] as follows. Let ${\mathscr{H}}$ be a gapped free fermion Hamiltonian such that ${\mathscr{T}}{\mathscr{H}}{\mathscr{T}}^{*}={\mathscr{H}}$, and let ${\mathscr{H}}(\phi)$ be the free fermion Hamiltonian obtained from ${\mathscr{H}}$ by inserting flux $\phi$. Then the spectrum of ${\mathscr{H}}(\phi)$ inherits the bulk spectrum of ${\mathscr{H}}$, but may in addition have eigenvalues in the bulk gap. As a function of $\phi$, these eigenvalues form bands within the gap of ${\mathscr{H}}$ corresponding to modes bound to the magnetic flux, see Figure 1. The time reversal symmetry forces the spectra of ${\mathscr{H}}(\phi)$ and ${\mathscr{H}}(-\phi)$ to be equal and for the time reversal invariant fluxes $\phi_{*}=0,\pm\pi$ the time reversal symmetry forces all eigenvalues of ${\mathscr{H}}(\phi_{*})$ to be doubly degenerate due to Kramers pairing. Energy bands satisfying these constraints come in two classes distinguished from each other as follows. Let $E_{F}$ be a generic Fermi energy in the bulk gap of the Hamiltonian ${\mathscr{H}}$ (for simplicity we assume that $E_{F}$ is not an eigenvalue of ${\mathscr{H}}(\phi_{*})$) and let $N\in{\mathbb{Z}}_{2}$ be the parity of the number of times a band crosses the Fermi energy in the interval $\phi\in[0,\pi)$. Then $N$ is invariant under any homotopy of the bands that respects the time reversal symmetry constraints. The Fu-Kane-Mele index of ${\mathscr{H}}$ (w.r.t. the chosen gap) is equal to $N$. The spectrum of Figure 1 corresponds to an ${\mathscr{H}}$ with a non-trivial Fu-Kane-Mele index. This characterisation of the Fu-Kane-Mele index parallels the characterisation for translation invariant free fermion systems through the bulk-boundary correspondence [27, 23, 12] where the bands associated to modes bound to the magnetic flux are replaced by bands formed by edge modes.
@@
Figure 1: The spectrum of a gapped free fermion Hamiltonian ${\mathscr{H}}(\phi)$ as a function of pierced flux. The filled grey regions are bulk spectrum. The bands represent spectrum associated to modes bound to the magnetic flux. The bands cross at time reversal invariant values of the flux $\phi=0,\pm\pi$ forming Kramers pairs represented by dots. The spectrum depicted here corresponds to a non-trivial insulator because for any Fermi energy $E_{F}$ in the bulk gap, the eigenvalue bands cross $E_{F}$ an odd number of times (in fact, ±$\pm 1$) in the interval $\phi\in[0,\pi)$.
@@
We now imagine preparing the system at $\phi=0$ in its ground state where all single particle states with energy below $E_{F}$ in the bulk gap are occupied, and increasing the flux $\phi$ adiabatically from 00 to $\pi$. The state is adiabatically transported along this path and at each value of $\phi\in[0,\pi]$ the occupied single particle states are indicated in the left panel of Figure 2. We see that one mode of the Kramers pair at $\phi=\pi$ is occupied while the other mode of the pair remains unoccupied. I instead the flux is decreased from $\phi=0$ to $\phi=-\pi$ we get a similar picture, except that now the other mode of the Kramers pair becomes occupied, see the right panel of Figure 2. The two states obtained from the ground state by either increasing to flux from 00 to $\pi$ or decreasing the flux from 00 to $-\pi$ are mapped into each other by time reversal.
@@
Under the assumption that $\omega$ is short-range entangled it follows that the states $\omega^{\pm}$ are unitarily equivalent (Proposition 3.8). The states $\omega^{\pm}$ therefore have unit vector representatives $\Omega^{\pm}\in{\mathcal{H}}$ in the GNS representation $(\pi,{\mathcal{H}})$ of either of them. Since they are time reversal images of each other, time reversal is implemented in the representation by an anti-unitary operator $T$ (Lemma 3.10), and $T^{2}$ implements fermion parity. It turns out that either $T^{2}\Omega^{+}=\Omega^{+}$ or $T^{2}\Omega^{+}=-\Omega^{+}$3.14). The first possibility correspond to a trivial index while the second possibility corresponds to a non-trivial ${\mathbb{Z}}_{2}$ index. In the non-trivial case we have that $\Omega^{-}=T\Omega^{+}$ and $\Omega^{+}$ form an orthogonal Kramers pair for $T$.
@@ 1.2 Organisation of the paper
The paper is organised as follows. In Section 2 we give precise definitions of the fermionic lattice systems and the class of pure states on these systems that we will study. We then state our main theorem: That there is a ${\mathbb{Z}}_{2}$-valued index associated to these states, that this index is stable under locally generated automorphisms which respect the symmetries, and that this index extends the Fu-Kane-Mele index for non-interacting quantum spin Hall effect. The rest of the paper is devoted to proving the main theorem. In Section 3 we define the index of a symmetric short-range entangled (SRE) state by carrying out the construction sketched above. In Section 4 we first prove that the index is well-defined in the sense that it is independent of various choices made in the definition. We then turn to its properties, namely multiplicativity under stacking and stability under locally generated automorphisms that respect the symmetries. We also show that symmetric product states have trivial index which allows us to extend the index to stably SRE states. Section 5 is devoted to the non-interacting situation. We first show that all translation-invariant, time reversal invariant quasi-free states are stably SRE and therefore have a well defined index. We then prove that our index agrees with the Fu-Kane-Mele index for translation invariant non-interacting spin Hall states. The next Section 6 collects these results to prove the main theorem. The appendices are devoted to some technical results. In Appendix A we prove a transitivity result for fermionic SRE states that is used in Section 3 to show that the ±$\pm\pi$ flux states are unitarily equivalent. In Appendix B we describe the free fermion analogue of the adiabatic flux insertion used in Section 5.
@@ 2.1 Observable algebras
for all sites $x,y\in{\mathbb{Z}}^{2}$ and all $i,j=1,\cdots,n$. For any $\Gamma\subset{\mathbb{Z}}^{2}$ we denote by ${\mathcal{A}}_{\Gamma}$ the unital $C^{*}$-subalgebra of ${\mathcal{A}}$ generated by the $\{a_{x,i}\}_{x\in\Gamma,i=1,\cdots,n}$. We also write ${\mathcal{A}}_{x}={\mathcal{A}}_{\{x\}}$ for any site $x\in{\mathbb{Z}}^{2}$. The fermion parity is the *-automorphism $\theta$ of ${\mathcal{A}}$ uniquely determined by $\theta(a_{x,i})=-a_{x,i}$ for all $x\in{\mathbb{Z}}^{2}$ and all $i=1,\cdots,n$. We have $\theta({\mathcal{A}}_{\Gamma})={\mathcal{A}}_{\Gamma}$ for all $\Gamma\subset{\mathbb{Z}}^{2}$. An operator $O\in{\mathcal{A}}$ is called even if $\theta(O)=O$, odd if $\theta(O)=-O$ and homogeneous if it is either even or odd. For any $\Gamma\subset{\mathbb{Z}}^{2}$ the subset of even operators of ${\mathcal{A}}_{\Gamma}$ is a $C^{*}$-subalgebra of ${\mathcal{A}}_{\Gamma}$ which we denote by ${\mathcal{A}}_{\Gamma}^{+}$. We will write $\Gamma\Subset{\mathbb{Z}}^{2}$ to mean that $\Gamma$ is a finite subset of ${\mathbb{Z}}^{2}$. If $A\in{\mathcal{A}}_{\Gamma}$ for some $\Gamma\ \Subset{\mathbb{Z}}^{2}$ then $A$ is said to be a local operator. The union of all local operators is a *-subalgebra of ${\mathcal{A}}$ which we denote by ${\mathcal{A}}^{\mathrm{loc}}$. It is dense in the topology of the C*-norm, see [14].
@@ 2.2 Locally generated automorphisms
see e.g. [9, 46].
@@ 2.3 Symmetries
We say that a TDI $F$ is time reversal symmetry preserving if $\tau(F_{x}(s))=-F_{x}(s)$ for all $s\in{\mathbb{R}}$ and all $x\in{\mathbb{Z}}^{2}$. Similarly, we say the TDI $F$ is $U(1)$ symmetry preserving if $\rho_{\phi}(F_{x}(s))=F_{x}(s)$ for all $\phi\in{\mathbb{R}}$, all $s\in{\mathbb{R}}$ and $x\in{\mathbb{Z}}^{2}$. A TDI that is simultaneously time reversal and $U(1)$ symmetry preserving is called a symmetry preserving TDI. If $F$ is a symmetry preserving TDI then the associated family of automorphisms $\alpha_{s}^{F}$ are called symmetry preserving LGAs, and satisfy $\rho_{\phi}\circ\alpha_{s}^{F}\circ\rho_{\phi}^{-1}=\tau\circ\alpha_{s}^{F}\circ\tau^{-1}=\alpha_{s}^{F}$.
@@ 2.4 Symmetric short-range entangled states
2.5 Examples : Non-interacting electrons with spin
We consider a single-particle Hilbert space ${\mathcal{K}}_{m}=l^{2}({\mathbb{Z}}^{2};{\mathbb{C}}^{2m})\simeq l^{2}({\mathbb{Z}}^{2})\otimes{\mathbb{C}}^{2m}$ for some positive integer $m$. We write the internal space as ${\mathbb{C}}^{2m}={\mathbb{C}}^{2}\otimes{\mathbb{C}}^{m}$ and fix an orthonormal tensor product basis $\{e_{\sigma,i}\}_{\sigma\in\{\uparrow,\downarrow\},i\in\{1,\cdots,m\}}$. We think of the label $\sigma$ as a spin-1/2 degree of freedom. Denote by $\{\delta_{x}\}_{x\in{\mathbb{Z}}^{2}}$ the orthonormal position basis of $l^{2}({\mathbb{Z}}^{2})$. Then the vectors $\{\delta_{x}\otimes e_{\sigma,i}\}$ form an orthonormal basis of ${\mathcal{K}}_{m}$. Denote by ${\mathscr{K}}$ the complex conjugation on ${\mathcal{K}}_{m}$ with respect to this basis, and let ${\mathscr{U}}_{T}$ be the unitary given by
@@ 2.5 Examples : Non-interacting electrons with spin
with respect to the tensor product decomposition ${\mathcal{K}}_{m}=l^{2}({\mathbb{Z}}^{2})\otimes{\mathbb{C}}^{2}\otimes{\mathbb{C}}^{m}$. We define an antiunitary time-reversal by
@@
Let ${\mathcal{A}}$ be the CAR algebra over ${\mathcal{K}}_{m}$. Then $\tau(a(f))=a(Tf)$ defines a fermionic time reversal on ${\mathcal{A}}$, i.e. $\tau^{2}=\theta$ and for any $\Gamma\subset{\mathbb{Z}}^{2}$ we have $\tau({\mathcal{A}}_{\Gamma})={\mathcal{A}}_{\Gamma}$, so $({\mathcal{A}},\tau)$ is a fermion system with time reversal.
Let ${\mathscr{P}}$ be an orthogonal projection on ${\mathcal{K}}_{m}$. The functional defined by
$\omega_{P}(a^{*}(f_{n^{\prime}})\cdots a^{*}(f_{1})a(g_{1})\cdots a(g_{n}))=\delta_{n^{\prime},n}\det\big{(}\langle g_{i},{\mathscr{P}}f_{j}\rangle_{i,j=1}^{n}\big{)}$ (2.11)
@@
2.5.1 The empty product state
Consider the trivial projection ${\mathscr{P}}=0$ on the single particle space ${\mathcal{K}}_{m}$. Then the corresponding quasi-free state $\omega_{\mathrm{empty}}=\omega_{{\mathscr{P}}}$ is characterized by $\omega_{\mathrm{empty}}(Q_{x})=0$ for all $x\in{\mathbb{Z}}^{2}$. It follows that $\omega_{\mathrm{empty}}$ is a product state because it restricts to each ${\mathcal{A}}_{x}$ as the unique state on ${\mathcal{A}}_{x}$ with zero expectation value for $Q_{x}$. Since ${\mathscr{T}}{\mathscr{P}}{\mathscr{T}}^{*}={\mathscr{P}}$ we see that $\omega_{\mathrm{empty}}$ is time reversal invariant, so $\omega_{\mathrm{empty}}$ is a symmetric product state.
2.5.2 Quasi-free translation invariant spin Hall states
Let us consider the single particle Hilbert space ${\mathcal{K}}_{2}=l^{2}({\mathbb{Z}}^{2};{\mathbb{C}}^{2}\otimes{\mathbb{C}}^{2})$ where the first ${\mathbb{C}}^{2}$ is the spin-1/2 degree of freedom, and the second ${\mathbb{C}}^{2}$ is some additional internal degree of freedom.
Using the Fourier transform we identify ${\mathcal{K}}_{2}=l^{2}({\mathbb{Z}}^{2};{\mathbb{C}}^{2}\otimes{\mathbb{C}}^{2})$ with $L^{2}({\mathbb{T}};{\mathbb{C}}^{2}\otimes{\mathbb{C}}^{2})$ where ${\mathbb{T}}$ is the Brillouin torus. Let us further identify $L^{2}({\mathbb{T}};{\mathbb{C}}^{2}\otimes{\mathbb{C}}^{2})\simeq L^{2}({\mathbb{T}};{\mathbb{C}}^{2})\oplus L^{2}({\mathbb{T}};{\mathbb{C}}^{2})$ where the first summand corresponds to $\sigma=\uparrow$ and the second summand corresponds to $\sigma=\downarrow$.
Let $\hat{\mathscr{P}}^{\uparrow}:{\mathbb{T}}\rightarrow\mathrm{Proj}({\mathbb{C}}^{2})$ be an analytic projection valued map of rank 1, which can be interpreted as a projection on $L^{2}({\mathbb{T}};{\mathbb{C}}^{2})$. Note that $E=\{(k,v)\in{\mathbb{T}}\times{\mathbb{C}}^{2}\,:\,v\in\mathrm{Ran}\hat{\mathscr{P}}^{\uparrow}(k)\}$ describes a line bundle over the torus. We define the Chern number $\mathrm{Ch}(\hat{\mathscr{P}}^{\uparrow})$ of $\hat{\mathscr{P}}^{\uparrow}$ as the Chern number of this line bundle. Denote by $\hat{\mathscr{P}}^{\downarrow}=\overline{\hat{\mathscr{P}}^{\uparrow}}$ the complex conjugate of $\hat{\mathscr{P}}^{\uparrow}$ using the conjugation ${\mathscr{K}}$. Then
$\hat{\mathscr{P}}_{SH}(k)=\hat{\mathscr{P}}^{\uparrow}(k)\oplus\hat{\mathscr{P}}^{\downarrow}(k)$ (2.12)
defines an analytic projection valued map of rank 2 which satisfies ${\mathscr{T}}\hat{\mathscr{P}}_{SH}(k){\mathscr{T}}^{*}=\hat{\mathscr{P}}_{SH}(-k)$ by construction. By Fourier transform we obtain the corresponding projection ${\mathscr{P}}_{SH}={\mathscr{P}}^{\uparrow}\oplus{\mathscr{P}}^{\downarrow}$ on ${\mathcal{K}}_{2}$ which satisfies ${\mathscr{T}}{\mathscr{P}}_{SH}{\mathscr{T}}^{*}={\mathscr{P}}_{SH}$. The corresponding quasi-free state $\omega_{SH}=\omega_{{\mathscr{P}}_{SH}}$ on ${\mathcal{A}}$ is time reversal invariant. We call any state of this form a quasi-free translation invariant spin Hall state. To avoid possible confusion here, we note that with this definition, a spin Hall state may be topologically trivial in the sense that it has both vanishing Chern number and trivial Fu-Kane-Mele ${\mathbb{Z}}_{2}$ index, see Section 5.
@@ 2.5.2 Quasi-free translation invariant spin Hall states
For later use we note that the analyticity of $\hat{\mathscr{P}}_{SH}$ implies that ${\mathscr{P}}_{SH}$ is exponentially local, i.e. there are constants $C,\eta>0$ such that
@@
$\left|\langle\delta_{x}\otimes e_{\sigma,i},\,{\mathscr{P}}_{SH}\,\delta_{y}\otimes e_{\sigma^{\prime},j}\rangle\right|\leq C{\mathrm{e}}^{-\eta\|x-y\|_{1}}$ (2.13)
@@
for all $x,y\in{\mathbb{Z}}^{2}$, all ${}\sigma,\sigma^{\prime}\in\{\uparrow\downarrow\}$, and all $i,j\in\{1,2\}$. We use the norm $\|x\|_{1}=\left|x_{1}\right|+\left|x_{2}\right|$ to measure spatial distances.
A typical example of the above setting is the Kane-Mele model without Rashba term and staggered chemical potential, see [27, Section 3.2]. In Section 5.1 we will show that
Lemma 2.3.
Quasi-free translation invariant spin Hall states are stably SRE.

### v2

@@ Abstract
We define a ${\mathbb{Z}}_{2}$-valued index for stably short-range entangled states of two-dimensional fermionic lattice systems with charge conservation and time reversal symmetry. The index takes its non-trivial value precisely if the ‘fluxon’, the state obtained by inserting a $\pi$-flux through the system, transforms under time reversal as part of a Kramers pair. This index extends the Fu-Kane-Mele index of free fermionic topological insulators to interacting systems.
Dedicated to the memory of Professor Huzihiro Araki
@@ 1 Introduction
The invariant that characterises the quantum spin Hall effect is the ${\mathbb{Z}}_{2}$-valued Fu-Kane-Mele index [33, 32, 26, 27]. It was argued in [48] [47] that this invariant has a physical interpretation in terms of time reversal transformation properties of defects bound to ±$\pm\pi$ magnetic fluxes. If the invariant is non-trivial then these defects form a Kramers pair for the time reversal symmetry. We review the argument below. While early works placed some importance on the role of spin, they all recognized that the invariant is in fact only relying on fermionic time reversal symmetry: no component of spin must be conserved for the invariant to be meaningful.
@@ 1.1 Idea behind the construction
The systems we consider are interacting generalisations of gapped free fermion Hamiltonians on the two-dimensional square lattice with a time reversal symmetry ${\mathscr{T}}$ that squares to $-\operatorname{\mathbbm{1}}$ (Altland-Zirnbauer class AII [3]). Such Hamiltonians come in two topological phases distinguished from each other by the ${\mathbb{Z}}_{2}$-valued Fu-Kane-Mele index. A nice characterisation of the index is obtained by piercing magnetic flux through the system [19] as follows. Let ${\mathscr{H}}$ be a gapped free fermion Hamiltonian such that ${\mathscr{T}}{\mathscr{H}}{\mathscr{T}}^{*}={\mathscr{H}}$, and let ${\mathscr{H}}(\phi)$ be the free fermion Hamiltonian obtained from ${\mathscr{H}}$ by inserting flux $\phi$. Then the spectrum of ${\mathscr{H}}(\phi)$ inherits the bulk spectrum of ${\mathscr{H}}$, but may in addition have eigenvalues in the bulk gap. As a function of $\phi$, these eigenvalues form bands within the gap of ${\mathscr{H}}$ corresponding to modes bound to the magnetic flux, see Figure 1. The time reversal symmetry forces the spectra of ${\mathscr{H}}(\phi)$ and ${\mathscr{H}}(-\phi)$ to be equal, and for the time reversal invariant fluxes $\phi_{*}=0,\pm\pi$ the time reversal symmetry forces all eigenvalues of ${\mathscr{H}}(\phi_{*})$ to be doubly degenerate due to Kramers pairing. Energy bands satisfying these constraints come in two classes distinguished from each other as follows. Let $E_{F}$ be a generic Fermi energy in the bulk gap of the Hamiltonian ${\mathscr{H}}$ (for simplicity we assume that $E_{F}$ is not an eigenvalue of ${\mathscr{H}}(\phi_{*})$) and let $N\in{\mathbb{Z}}_{2}$ be the parity of the number of times a band crosses the Fermi energy in the interval $\phi\in[0,\pi)$. Then $N$ is invariant under any homotopy of the bands that respects the time reversal symmetry constraints. The Fu-Kane-Mele index of ${\mathscr{H}}$ (w.r.t. the chosen gap) is equal to $N$. The spectrum of Figure 1 corresponds to an ${\mathscr{H}}$ with a non-trivial Fu-Kane-Mele index. This characterisation of the Fu-Kane-Mele index parallels the characterisation for translation invariant free fermion systems through the bulk-boundary correspondence [28, 24, 15] where the bands associated to modes bound to the magnetic flux are replaced by bands formed by edge modes.
@@
Figure 1: The spectrum of a gapped free fermion Hamiltonian ${\mathscr{H}}(\phi)$ as a function of pierced flux. The filled grey regions are bulk spectrum. The bands represent spectrum associated to modes bound to the magnetic flux. The bands cross at time reversal invariant values of the flux $\phi_{*}=0,\pm\pi$ forming Kramers pairs represented by dots. The spectrum depicted here corresponds to a non-trivial insulator because for any Fermi energy $E_{F}$ in the bulk gap, the eigenvalue bands cross $E_{F}$ an odd number of times (in fact, ±$\pm 1$) in the interval $\phi\in[0,\pi)$.
@@
We now imagine preparing the system at $\phi=0$ in its ground state where all single particle states with energy below $E_{F}$ in the bulk gap are occupied, and increasing the flux $\phi$ adiabatically from 00 to $\pi$. The state is adiabatically transported along this path and at each value of $\phi\in[0,\pi]$ the occupied single particle states are indicated in the left panel of Figure 2. We see that one mode of the Kramers pair at $\phi=\pi$ is occupied while the other mode of the pair remains unoccupied. If the flux is instead decreased from $\phi=0$ to $\phi=-\pi$ we get a similar picture, except that now the other mode of the Kramers pair becomes occupied, see the right panel of Figure 2. The two states obtained from the ground state by either increasing the flux from 00 to $\pi$ or decreasing the flux from 00 to $-\pi$ are mapped into each other by time reversal.
@@
Under the assumption that $\omega$ is short-range entangled it follows that the states $\omega^{\pm}$ are unitarily equivalent (Proposition 3.8). The states $\omega^{\pm}$ therefore have unit vector representatives $\Omega^{\pm}\in{\mathcal{H}}$ in the GNS representation $(\pi,{\mathcal{H}})$ of either of them. Since they are time reversal images of each other, time reversal is implemented in the representation by an antiunitary operator $T$ (Lemma 3.10). It turns out that either $T^{2}\Omega^{+}=\Omega^{+}$ or $T^{2}\Omega^{+}=-\Omega^{+}$ (Lemma 3.14). The first possibility corresponds to a trivial value for the many-body index, while the second possibility corresponds to a non-trivial value of the index. In the non-trivial case we have that $\Omega^{-}=T\Omega^{+}$ and $\Omega^{+}$ form an orthogonal Kramers pair for $T$.
@@ 1.2 Organisation of the paper
[…]
@@ 2.1 Observable algebras
for all sites $x,y\in{\mathbb{Z}}^{2}$ and all $i,j=1,\cdots,n$. For any $\Gamma\subset{\mathbb{Z}}^{2}$ we denote by ${\mathcal{A}}_{\Gamma}$ the unital $C^{*}$-subalgebra of ${\mathcal{A}}$ generated by the $\{a_{x,i}\}_{x\in\Gamma,i=1,\cdots,n}$. We also write ${\mathcal{A}}_{x}={\mathcal{A}}_{\{x\}}$ for any site $x\in{\mathbb{Z}}^{2}$. The fermion parity is the *-automorphism $\theta$ of ${\mathcal{A}}$ uniquely determined by $\theta(a_{x,i})=-a_{x,i}$ for all $x\in{\mathbb{Z}}^{2}$ and all $i=1,\ldots,n$. We have $\theta({\mathcal{A}}_{\Gamma})={\mathcal{A}}_{\Gamma}$ for all $\Gamma\subset{\mathbb{Z}}^{2}$. An operator $O\in{\mathcal{A}}$ is called even if $\theta(O)=O$, odd if $\theta(O)=-O$ and homogeneous if it is either even or odd. For any $\Gamma\subset{\mathbb{Z}}^{2}$ the subset of even operators of ${\mathcal{A}}_{\Gamma}$ is a $C^{*}$-subalgebra of ${\mathcal{A}}_{\Gamma}$ which we denote by ${\mathcal{A}}_{\Gamma}^{+}$. We will write $\Gamma\Subset{\mathbb{Z}}^{2}$ to mean that $\Gamma$ is a finite subset of ${\mathbb{Z}}^{2}$. If $A\in{\mathcal{A}}_{\Gamma}$ for some $\Gamma\ \Subset{\mathbb{Z}}^{2}$ then $A$ is said to be a local operator. The union of all local operators is a *-subalgebra of ${\mathcal{A}}$ which we denote by ${\mathcal{A}}^{\mathrm{loc}}$. It is dense in the topology of the C*-norm, see [17].
@@ 2.2 Locally generated automorphisms
see e.g. [12, 46]. Since $\alpha^{F}_{s}$ is bounded, it extends to an automorphism of the whole algebra ${\mathcal{A}}$, which we denote by the same symbol.
@@ 2.3 Symmetries
We say that a TDI $F$ is time reversal symmetry preserving if $\tau(F_{x}(s))=-F_{x}(s)$ for all $s\in{\mathbb{R}}$ and all $x\in{\mathbb{Z}}^{2}$. Similarly, we say the TDI $F$ is $U(1)$ symmetry preserving if $\rho_{\phi}(F_{x}(s))=F_{x}(s)$ for all $\phi\in{\mathbb{R}}$, all $s\in{\mathbb{R}}$ and $x\in{\mathbb{Z}}^{2}$. A TDI that is simultaneously time reversal and $U(1)$ symmetry preserving is called a symmetry preserving TDI. If $F$ is a symmetry preserving TDI then the associated family of automorphisms $\alpha_{s}^{F}$ satisfy
$\rho_{\phi}\circ\alpha_{s}^{F}\circ\rho_{\phi}^{-1}=\alpha_{s}^{F}=\tau\circ\alpha_{s}^{F}\circ\tau^{-1}$ (2.9)
and $\alpha_{s}^{F}$ are called symmetry preserving LGAs.
@@ 2.4 Symmetric short-range entangled states
2.5 Examples : Quasi-free symmetric states and the Fu-Kane-Mele index
We consider a single-particle Hilbert space ${\mathcal{K}}_{m}=l^{2}({\mathbb{Z}}^{2};{\mathbb{C}}^{2m})\simeq l^{2}({\mathbb{Z}}^{2})\otimes{\mathbb{C}}^{2m}$ for some positive integer $m$. We write the internal space as ${\mathbb{C}}^{2m}={\mathbb{C}}^{2}\otimes{\mathbb{C}}^{m}$ and fix an orthonormal tensor product basis $\{e_{\sigma,i}\}_{\sigma\in\{\uparrow,\downarrow\},i\in\{1,\ldots,m\}}$. Although this will not play a role in the following, one may think of the label $\sigma$ as a spin-1/2 degree of freedom. Denote by $\{\delta_{x}\}_{x\in{\mathbb{Z}}^{2}}$ the orthonormal position basis of $l^{2}({\mathbb{Z}}^{2})$. Then the vectors $\{\delta_{x}\otimes e_{\sigma,i}\}$ form an orthonormal basis of ${\mathcal{K}}_{m}$. Denote by ${\mathscr{K}}$ the complex conjugation on ${\mathcal{K}}_{m}$ with respect to this basis, and let ${\mathscr{U}}_{T}$ be the unitary given by
@@ 2.5 Examples : Quasi-free symmetric states and the Fu-Kane-Mele index
with respect to the tensor product decomposition ${\mathcal{K}}_{m}=l^{2}({\mathbb{Z}}^{2})\otimes{\mathbb{C}}^{2}\otimes{\mathbb{C}}^{m}$. We define an antiunitary time reversal on ${\mathcal{K}}_{m}$ by
@@
Let ${\mathcal{A}}$ be the CAR algebra over ${\mathcal{K}}_{m}$. Then $\tau(a(f))=a({\mathscr{T}}f)$ defines a fermionic time reversal on ${\mathcal{A}}$, i.e. $\tau^{2}=\theta$ and for any $\Gamma\subset{\mathbb{Z}}^{2}$ we have $\tau({\mathcal{A}}_{\Gamma})={\mathcal{A}}_{\Gamma}$, so $({\mathcal{A}},\tau)$ is a fermion system with time reversal.
If ${\mathscr{P}}$ is an orthogonal projector on ${\mathcal{K}}_{m}$ then the functional $\omega_{{\mathscr{P}}}$ defined by
$\omega_{{\mathscr{P}}}(a^{*}(f_{n^{\prime}})\cdots a^{*}(f_{1})a(g_{1})\cdots a(g_{n}))=\delta_{n^{\prime},n}\det\big{(}\langle g_{i},{\mathscr{P}}f_{j}\rangle_{i,j=1}^{n}\big{)}$ (2.15)
@@
A projector ${\mathscr{P}}$ on ${\mathcal{K}}_{m}$ is exponentially local if there are constants $C,\eta>0$ such that
@@
$\left|\langle\delta_{x}\otimes e_{\sigma,i},\,{\mathscr{P}}\,\delta_{y}\otimes e_{\sigma^{\prime},j}\rangle\right|\leq C{\mathrm{e}}^{-\eta\|x-y\|_{1}}$ (2.16)
@@
for all $x,y\in{\mathbb{Z}}^{2}$, all ${}\sigma,\sigma^{\prime}\in\{\uparrow\downarrow\}$, and all $i,j\in\{1,\ldots,m\}$. We use the norm $\|x\|_{1}=\left|x_{1}\right|+\left|x_{2}\right|$ to measure spatial distances on ${\mathbb{Z}}^{2}$.
A projector ${\mathscr{P}}$ on ${\mathcal{K}}_{m}$ is time reversal invariant if ${\mathscr{T}}{\mathscr{P}}{\mathscr{T}}^{*}={\mathscr{P}}$. If ${\mathscr{P}}$ is also exponentially local then we can regard ${\mathscr{P}}$ as the Fermi projection of a gapped exponentially local and time reversal invariant Hamiltonian ${\mathscr{H}}$ at some Fermi energy $\mu$. For concreteness, we may take ${\mathscr{H}}=\operatorname{\mathbbm{1}}-{\mathscr{P}}$ and $\mu=1/2$. Let us define the corresponding family of flux Hamiltonians ${\mathscr{H}}_{\phi}$ by
${\mathscr{H}}_{\phi}(x,i;y,j)=\begin{cases}{\mathrm{e}}^{{\mathrm{i}}\phi\,\mathrm{sgn}(x_{2}-y_{2})}{\mathscr{H}}(x,i;y,j)&\text{if }x_{1},y_{1}\leq 0\\ {\mathscr{H}}(x,i;y,j)&\text{otherwise}\end{cases}$ (2.17)
where for any operator ${\mathscr{A}}\in{\mathcal{B}}({\mathscr{K}}_{m})$ we write ${\mathscr{A}}(x,i;y,j)=\langle\delta_{x}\otimes e_{i},{\mathscr{A}}\,\delta_{y}\otimes e_{j}\rangle$ for its matrix elements w.r.t. the basis $\{\delta_{x}\otimes e_{i}\}_{x,i}$ (to simplify notations, we shall from here on drop the additional index $\sigma$ whenever it is irrelevant). Note that $\phi\mapsto{\mathscr{H}}_{\phi}$ is $2\pi$-periodic.
The Fu-Kane-Mele index of ${\mathscr{P}}$ is given by a spectral flow [19]
$\operatorname{\mathrm{ind}_{\mathrm{FKM}}}({\mathscr{P}}):=\operatorname{\mathrm{SF}}_{\mu}([0,\pi]\ni\phi\mapsto{\mathscr{H}}_{\phi})\,\mod 2$ (2.18)
for any $\mu\in(0,1)$ in the gap of ${\mathscr{H}}$.
A quasi-free state $\omega_{{\mathscr{P}}}$ corresponding to an exponentially local and time reversal invariant projector ${\mathscr{P}}$ is called an $\operatorname{\mathrm{AII}}$ state, referring to the relevant Altland-Zirnbauer symmetry class.
A special case of $\operatorname{\mathrm{AII}}$ states is obtained by taking the trivial projection ${\mathscr{P}}=0$ on the single particle space ${\mathcal{K}}_{m}$. Then the corresponding quasi-free state $\omega_{\mathrm{empty}}=\omega_{{\mathscr{P}}}$ is characterized by $\omega_{\mathrm{empty}}(Q_{x})=0$ for all $x\in{\mathbb{Z}}^{2}$. It follows that $\omega_{\mathrm{empty}}$ is a product state because it restricts to each ${\mathcal{A}}_{x}$ as the unique state on ${\mathcal{A}}_{x}$ with zero expectation value for $Q_{x}$. Since ${\mathscr{T}}{\mathscr{P}}{\mathscr{T}}^{*}={\mathscr{P}}$ we see that $\omega_{\mathrm{empty}}$ is time reversal invariant, so $\omega_{\mathrm{empty}}$ is a symmetric product state, in particular SRE.